OpenStudy (anonymous):

carla paid 462.25 for textbooks with 7.5% sales tax included what was the price of the text books

OpenStudy (anonymous):

thanks

OpenStudy (zarkon):

430

OpenStudy (anonymous):

What's up! 462.25 = x*1.075 x = 462.25/1.075